# FC–PINO: High-Precision Physics-Informed Neural Operators via Fourier Continuation

### v2

@@ 2.2.3 Other Physics-informed Approaches
Various approaches enforce physics laws in surrogate models without using physics losses. This can be achieved, for instance, by using projection layers [45; 46; 47] or by finding optimal linear combinations of learned basis functions that solve a PDE-constrained optimization problem [48; 49]. Another approach is to leverage well-studied inherent properties of the solution operator. For instance, incompressibility $\nabla\cdot u=0$ can be enforced by representing $u$ as the curl of a vector field $v$, $u=\nabla\times v$, and learning $v$ directly [50], which guarantees the resulting $u$ is divergence-free. As another example, machine learning surrogates can preserve the symplectic structure underlying Hamiltonian systems [51; 52; 53; 54; 55], to ensure conservation of energy and other fundamental invariants, which leads to predictions that remain physically realistic over long time horizons.
@@
Hybrid strategies are also possible, where some constraints can be directly enforced in the architecture while others are imposed via physics losses during training. For example, in incompressible Navier–Stokes problems, a physics loss can penalize deviations from the momentum equation, while incompressibility can be explicitly enforced using projections as demonstrated in [45; 46].
@@ 2.3 Computing Derivatives for Physics Losses
However, finite differences face the same challenges as the corresponding numerical solvers: they can require high-resolution grids to achieve accurate results. If the grid is too coarse, derivative approximations can be poor, leading to significant errors. Consequently, finite differences can become computationally expensive or even intractable for multi-scale systems or for dynamics that vary rapidly in space or time.
@@
 Derivatives can be evaluated pointwise using automatic differentiation, which applies the chain rule to the sequence of operations in the model. Tools like Autograd [56] automate this by constructing a computational graph during the forward pass and using reverse-mode differentiation to compute gradients in the backward pass. Automatic differentiation is often preferred over finite differences in PINNs [19] because (1) it yields extremely accurate derivatives even at low resolutions, (2) it requires only a single function evaluation, and (3) it enables efficient computation of higher-order derivatives with minimal overhead. In contrast, finite differences require more evaluations and are more prone to numerical errors.
However, automatic differentiation has drawbacks: all operations must be differentiable, and storing intermediate values in the computational graph can lead to high memory consumption in deep models. It may also be slower than finite differences for problems where the latter are sufficiently accurate on low-resolution grids. Moreover, deep compositions in the physics loss can exacerbate issues such as vanishing or exploding gradients during backpropagation.
@@
 Spectral differentiation (or Fourier differentiation) is a fast and memory-efficient method to approximate derivatives by working in frequency space [57; 58]. Spectral differentiation is especially appealing when the solution is smooth and periodic, as it delivers very high accuracy, even at relatively low and moderate resolution, and can be implemented efficiently via the fast Fourier transform (FFT). Spectral differentiation exploits the fact that any sufficiently smooth and periodic function $f:[0,2\pi]\to\mathbb{R}$ can be represented by its Fourier series
$f(x)\;=\;\sum_{k=-\infty}^{\infty}\widehat{f}_{k}\,e^{ikx},\quad\widehat{f}_{k}\;=\;\frac{1}{2\pi}\int_{0}^{2\pi}f(x)e^{-ikx}\,\mathrm{d}x.$
Taking derivatives in physical space corresponds to simple multiplications by powers of $ik$ in frequency space, where $k$ represents the wavenumber in Fourier space. More precisely, let $n$ be the number of equidistant grid points. Then, the $m$-th derivative of $f$ has Fourier coefficients
@@
Then, after applying the inverse FFT, we recover the $m$-th derivative $\partial^{m}f(2\pi j/n)$ of $f$ at the $n$ equidistant grid points.
A major advantage of spectral differentiation is that higher derivatives are obtained simply via multiplication by $(ik)^{m}$ for the same computational cost irrespective of the derivative order, whereas finite differences require wider stencils and greater computational cost as the derivative order increases. Spectral differentiation is also much more memory-efficient than automatic differentiation. When applied to high-dimensional PDEs and deep networks, automatic differentiation must record the entire computational graph and intermediate values, often leading to prohibitive memory overhead, complex implementation, and slow backward passes when training with physics losses. In contrast, spectral differentiation requires only the storage of the function values and its Fourier coefficients.
However, like other spectral techniques, spectral differentiation is inherently tailored to periodic functions. If the target function is non-periodic, the resulting Gibbs phenomenon and poor convergence of the Fourier series can degrade derivative accuracy (see Figure˜3), potentially undermining the reliability of physics losses for training models [57; 58]. This highlights the need for modified spectral differentiation methods that can robustly handle non-periodic functions for physics-informed training of machine learning models.
@@ 3 Methods
A central challenge in spectral differentiation in PINO is that Fourier series are well-defined only for periodic functions [57]. For non-periodic problems, one must first construct a periodic extension over a larger domain, commonly by padding, where the model produces outputs on an extended interval (e.g., [1,2][-1,2]), while losses are evaluated solely on the original domain (e.g., $[0,1]$). This partial observability where the model is optimized using data only from a subinterval becomes particularly problematic in Fourier layers, which rely on the discrete Fourier transform (DFT). The DFT can be written explicitly as a Vandermonde matrix, with entries $F_{jk}=e^{-2\pi ijk/N}$, mapping signals to their frequency coefficients. Then, the condition number of this Vandermonde matrix (i.e., ratio of its largest to smallest singular value) grows rapidly with resolution [20; 21; 22]. As a result, the transform becomes ill-conditioned, and small perturbations in the input can produce disproportionately large deviations in the output. This is exacerbated when the function being transformed is non-periodic or exhibits discontinuous derivatives at the boundaries due to padding, in which case spurious high-frequency components are introduced. Such ill-conditioning not only degrades solution accuracy but also poses a significant challenge for gradient-based optimization in PINO models, underscoring the need for careful boundary treatment and continuation methods.
@@ 3.2.1 Motivation
Ideally, to apply Fourier methods to problems with non-periodic solutions, one must convert the signal of interest to a periodic function in a well-conditioned manner. This is precisely the purpose of Fourier continuation (FC) methods. They extend non-periodic signals to periodic ones over an extended domain through polynomial interpolation, while achieving empirically superalgebraic convergence rates. We will introduce two Fourier continuation methods that can be integrated into FC–PINO: FC–Legendre and (Accelerated) FC–Gram [26]. Figure˜1 displays examples of periodic extensions using FC–Legendre and FC–Gram on 1D and 2D non-periodic signals.
@@ 3.2.2 General Setting
For functions in higher dimensions, the one-dimensional extension is applied along each direction iteratively using the same extension matrices, i.e. first along the first dimension (left/right), then along the second dimension (top/bottom), and so on.
@@ 3.2.4 FC–Gram
We now describe the (Accelerated) FC–Gram method from [26] which is inspired from earlier FC–Gram methods [59; 23]. See [26] for more details and illustrations of FC–Gram.
@@
Tuning FC–Gram is more involved, owing to its localized structure and the multiple construction stages, each with several parameters that can be adjusted. However, after extensive tuning conducted in previous FC–Gram works [59; 23; 26], fixed values of most parameters have been determined to provide an appropriate trade-off between overall computational cost and stability which works well in most settings. Then, we can simply focus on tuning the polynomial degree $d$, and refining the resolution as necessary since increasing the signal resolution can help further reduce PDE residuals. Higher $d$ allows to better resolve finer boundary features and higher-order derivatives, while excessively high values can cause ill-conditioning, noise overfitting, or oscillatory artifacts.
Overall, FC–Gram combines flexibility and robustness: its localized blend-to-zero approach produces smooth, stable derivative fields, and even with minimal tuning, it consistently achieves leads to high-precision solutions when paired with FC–PINO across a wide range of problems. More details about tuning FC–Gram for high-precision applications are provided in Section˜B.2.2.
@@ 3.3 Comparison of FC–Legendre and FC–Gram
FC–Gram and FC–Legendre employ fundamentally different continuation strategies, each with unique pros and cons. As illustrated in Figure˜1, FC–Legendre constructs the extension by fitting a single Legendre polynomial of degree (2d1)(2d-1) spanning the entire extended interval between the left and right boundaries. In contrast, FC–Gram takes a more local approach, projecting boundary data onto lower-order polynomial bases of degree $d$ separately on each boundary and then smoothly blending the extensions to zero outside the domain. While both approaches yield good results with minimal tuning alongside FC–PINO, one can offer advantages over the other in high-precision settings.
@@
In contrast, FC–Gram excels for signals that are less smooth, noisy, or contain discontinuities, as it provides greater robustness and effective localized error control. By constructing localized orthonormal polynomial bases on each boundary and smoothly blending the extension to zero, it confines boundary errors and high-frequency noise within the extension region, preventing their propagation into the main domain. This approach suppresses spurious oscillations near boundaries, resulting in cleaner derivatives without the need for manual filtering. Although FC–Gram is more computationally intensive and requires higher precision arithmetic, its QR factorizations performed on smaller local matrices produce better-conditioned continuation operators, which improve numerical stability and reduce sensitivity to round-off errors. Overall, its enhanced robustness and localized error confinement makes it especially well-suited for noisy or non-smooth problems.
@@ 3.4 Error Convergence for Solutions and Derivatives in FC–PINO
For FC Gram, the continuation error on the physical interval consists of two parts, as demonstrated in [60; 61]. The dominant term is the interpolation error from matching $d$ boundary values, which converges at order $\mathcal{O}(N^{-d})$. In addition, there is a blending term caused by forcing the continuation to vanish smoothly outside the interval. This blending error does not formally converge with $N$, but can be made numerically negligible. When computing a derivative of order $k$, the interpolation error is reduced to order $\mathcal{O}(N^{-(d-k)})$, while the blending error remains bounded and does not affect the overall convergence rate.
@@
Now, FC–PINO is based on the Fourier neural operator, which enjoys universal approximation for sufficiently smooth operator mappings [5], i.e., one can make the approximation error $\|u_{\theta}-\mathcal{G}(f)\|$ arbitrarily small. By the Sobolev embedding theorem, with sufficient regularity, the error in the derivatives can also be controlled and kept arbitrarily small.
@@ 4.1.1 Problem Setting
[…]
@@ 4.1.3 Learning the Family of Solutions
It is known that for the Burgers equation to have a smooth self-similar solution, $\lambda$ has to take the form $\lambda=1/(2i+2)$ [30]. Thus we obtain a collection of values of $\lambda=1/(2i+2)$ by randomly sampling $i\in[0,20]$. By training the neural operator on the corresponding Burgers’ equations, we effectively learn the solution operator over arbitrary values of $\lambda$ corresponding to smooth self-similar solutions.
@@
We first pretrain the proposed FC–PINO architecture with FC–Gram for $3000$ epochs across the selected subset of values of $\lambda$. Then, we fine-tune the pretrained model on individual values of $\lambda$, for $\lambda\in\{1/2,1/12,1/22,1/32,1/52\}$, using the second-order L–BFGS optimizer for 40,00040,\!000 epochs. We embed $\lambda$ in the architecture by representing it as the amplitude of sines and cosines in a NeRF-style embedding [62], as
@@
for some integer $K>0$, where $x$ denotes the discretized spatial coordinates. FNOs naturally leverage sines and cosines through the Fourier transform to learn mappings by capturing interactions between frequency components. Using an encoding with a richer frequency spectrum than a constant $\lambda$ function can allow more efficient learning of complex spectral dynamics. In this numerical experiment, $K=2$ proved sufficient.

### v3

@@ 2.2.3 Other Physics-informed Approaches
Various approaches enforce physics laws in surrogate models without using physics losses. This can be achieved, for instance, by using projection layers (Jiang et al., 2020; Duruisseaux et al., 2024; Harder et al., 2024) or by finding optimal linear combinations of learned basis functions that solve a PDE-constrained optimization problem (Negiar et al., 2022; Chalapathi et al., 2024). Another approach is to leverage well-studied inherent properties of the solution operator. For instance, incompressibility $\nabla\cdot u=0$ can be enforced by representing $u$ as the curl of a vector field $v$, $u=\nabla\times v$, and learning $v$ directly (Mohan et al., 2023), which guarantees the resulting $u$ is divergence-free. As another example, machine learning surrogates can preserve the symplectic structure underlying Hamiltonian systems (Burby et al., 2020; Jin et al., 2020; Duruisseaux et al., 2023a, b; Valperga et al., 2022), to ensure conservation of energy and other fundamental invariants, which leads to predictions that remain physically realistic over long time horizons.
@@
Hybrid strategies are also possible, where some constraints can be directly enforced in the architecture while others are imposed via physics losses during training. For example, in incompressible Navier–Stokes problems, a physics loss can penalize deviations from the momentum equation, while incompressibility can be explicitly enforced using projections as demonstrated in (Jiang et al., 2020; Duruisseaux et al., 2024).
@@ 2.3 Computing Derivatives for Physics Losses
However, finite differences face the same challenges as the corresponding numerical solvers: they can require high-resolution grids to achieve accurate results. If the grid is too coarse, derivative approximations can be poor, leading to significant errors. Consequently, finite differences can become computationally expensive or even intractable for multi-scale systems or for dynamics that vary rapidly in space or time. Note that higher-order finite difference schemes or stencils can sometimes help mitigate these issues.
@@
Derivatives can be evaluated pointwise using automatic differentiation, which applies the chain rule to the sequence of operations in the model. Tools like Autograd (Maclaurin et al., 2015) automate this by constructing a computational graph during the forward pass and using reverse-mode differentiation to compute gradients in the backward pass. Automatic differentiation is often preferred over finite differences in PINNs (Baydin et al., 2017) because (1) it yields extremely accurate derivatives even at low resolutions, (2) it requires only a single function evaluation, and (3) it enables efficient computation of higher-order derivatives with minimal overhead. In contrast, finite differences require more evaluations and are more prone to numerical errors.
However, automatic differentiation has drawbacks: all operations must be differentiable, and storing intermediate values in the computational graph can lead to high memory consumption in deep models. It may also be slower than finite differences for problems where the latter are sufficiently accurate on low-resolution grids. Moreover, deep compositions in the physics loss can exacerbate vanishing or exploding gradient issues during backpropagation.
@@
Spectral differentiation (or Fourier differentiation) is a fast and memory-efficient method to approximate derivatives by working in frequency space (Trefethen, 2000; Boyd, 2013). Spectral differentiation is especially appealing when the solution is smooth and periodic, as it delivers very high accuracy, even at relatively low and moderate resolution, and can be implemented efficiently via the fast Fourier transform (FFT). Spectral differentiation exploits the fact that any sufficiently smooth and periodic function $f:[0,2\pi]\to\mathbb{R}$ can be represented by its Fourier series
$f(x)\;=\;\sum_{k=-\infty}^{\infty}\widehat{f}_{k}\,e^{ikx},\qquad\widehat{f}_{k}\;=\;\frac{1}{2\pi}\int_{0}^{2\pi}f(x)e^{-ikx}\,\mathrm{d}x.$
Taking derivatives in physical space corresponds to simple multiplications by powers of $ik$ in frequency space, where $k$ represents the wavenumber. More precisely, let $n$ be the number of equidistant grid points. Then, the $m$-th derivative of $f$ has Fourier coefficients
@@
After applying an inverse FFT, we recover the $m$-th derivative $\partial^{m}f(2\pi j/n)$ of $f$ at the $n$ equidistant grid points. A major advantage of spectral differentiation is that higher derivatives are obtained via multiplication by $(ik)^{m}$ for the same computational cost irrespective of the derivative order, whereas finite differences require wider stencils and greater computational cost as the derivative order increases.
Spectral differentiation is also much more memory‐efficient than automatic differentiation. When applied to high‐dimensional PDEs and deep networks, automatic differentiation must record the entire computational graph and intermediate values, often leading to prohibitive memory overhead, complex implementation, and slow backward passes when training with physics losses. In contrast, spectral differentiation requires only the storage of the function values and its Fourier coefficients.
Note that spectral differentiation is highly efficient primarily when the underlying function is noise-free. When the data is noisy, it can become ill-conditioned and lead to a significant loss of accuracy. To mitigate the amplification of errors associated with high-frequency noise, spectral differentiation is often combined with low-pass spectral filtering in noisy settings, which reduces the impact of noise but inevitably degrades the accuracy of the resulting spectral derivatives. In these situations, alternative differentiation techniques, such as high-order finite difference schemes, can sometimes yield more reliable and accurate results.
Unfortunately, like other spectral techniques, spectral differentiation is inherently tailored to periodic functions. If the target function is non-periodic, the resulting Gibbs phenomenon and poor convergence of the Fourier series can degrade derivative accuracy (see Figure˜3), potentially undermining the reliability of physics losses for training models (Trefethen, 2000; Boyd, 2013). This highlights the need for modified spectral differentiation methods that can robustly handle non-periodic functions for physics-informed training of machine learning models.
@@ 3 Methods
A central challenge in spectral differentiation in PINO is that Fourier series are well-defined only for periodic functions (Trefethen, 2000). For non-periodic problems, one must first construct a periodic extension over a larger domain, commonly by padding, where the model produces outputs on an extended interval (e.g., [1,2][-1,2]), while losses are evaluated solely on the original domain (e.g., $[0,1]$). This partial observability where the model is optimized using data only from a subinterval becomes particularly problematic in Fourier layers, which rely on the discrete Fourier transform (DFT). The DFT can be written explicitly as a Vandermonde matrix, with entries $F_{jk}=e^{-2\pi ijk/N}$, mapping signals to their frequency coefficients. Then, the condition number of this Vandermonde matrix (i.e., ratio of its largest to smallest singular value) grows rapidly with resolution (Higham, 2002; Press et al., 2007; Adcock et al., 2014). As a result, the transform becomes ill-conditioned, and small perturbations in the input can produce disproportionately large deviations in the output. This is exacerbated when the function being transformed is non-periodic or exhibits discontinuous derivatives at the boundaries due to padding, in which case spurious high-frequency components are introduced. Such ill-conditioning not only degrades solution accuracy but also poses a significant challenge for gradient-based optimization in PINO models, underscoring the need for careful boundary treatment and continuation methods.
@@ 3.2.1 Motivation
Ideally, to apply Fourier methods to problems with non-periodic solutions, one must convert the signal of interest to a periodic function in a well-conditioned manner. This is precisely the purpose of Fourier continuation (FC) methods. They extend non-periodic signals to periodic ones over an extended domain through polynomial interpolation, while achieving empirically superalgebraic convergence rates. We will introduce two Fourier continuation methods that can be integrated into FC–PINO: FC–Legendre and (Accelerated) FC–Gram (Amlani and Bruno, 2016). Figure˜1 displays examples of periodic extensions using FC–Legendre and FC–Gram on 1D and 2D non-periodic signals.
@@ 3.2.2 General Setting
For functions in higher dimensions, the one‐dimensional extension is applied along each direction iteratively using the same extension matrices, i.e. first along the first dimension (left/right), then along the second dimension (top/bottom), and so on.
@@ 3.2.4 FC–Gram
We now describe the (Accelerated) FC–Gram method from (Amlani and Bruno, 2016) which is inspired from earlier FC–Gram methods (Bruno et al., 2007; Albin and Bruno, 2011). See (Amlani and Bruno, 2016) for more details and illustrations of FC–Gram.
@@
Tuning FC–Gram is more involved, owing to its localized structure and the multiple construction stages, each with several parameters that can be adjusted. However, after extensive tuning conducted in previous FC–Gram works (Bruno et al., 2007; Albin and Bruno, 2011; Amlani and Bruno, 2016), fixed values of most parameters have been determined to provide an appropriate trade-off between overall computational cost and stability which works well in most settings. Then, we can simply focus on tuning the polynomial degree $d$, and refining the resolution as necessary since increasing the signal resolution can help further reduce PDE residuals. Higher $d$ allows to better resolve finer boundary features and higher-order derivatives, while excessively high values can cause ill-conditioning, noise overfitting, or oscillatory artifacts.
Overall, FC–Gram combines flexibility and robustness: its localized blending approach produces smooth, stable derivative fields, and even with minimal tuning, it consistently achieves leads to high-precision solutions when paired with FC–PINO across a wide range of problems. More details about tuning FC–Gram for high-precision applications are provided in Section˜B.2.2.
@@ 3.3 Comparison of FC–Legendre and FC–Gram
FC–Gram and FC–Legendre employ fundamentally different continuation strategies, each with unique pros and cons. As illustrated in Figure˜1, FC–Legendre constructs the extension by fitting a single Legendre polynomial of degree (2d1)(2d-1) spanning the entire extended interval between the left and right boundaries. In contrast, FC–Gram takes a more local approach, projecting boundary data onto lower-order polynomial bases of degree $d$ separately on each boundary and then smoothly blending the extensions outside the domain. While both approaches yield good results with minimal tuning alongside FC–PINO, one can offer advantages over the other in high-precision settings.
@@
In contrast, FC–Gram excels for signals that are less smooth, noisy, or contain discontinuities, as it provides greater robustness and effective localized error control. By constructing localized orthonormal polynomial bases on each boundary and smoothly blending the extension, it confines boundary errors and high-frequency noise within the extension region, preventing their propagation into the main domain. This approach suppresses spurious oscillations near boundaries, resulting in cleaner derivatives without the need for manual filtering. Although FC–Gram is more computationally intensive and requires higher precision arithmetic, its QR factorizations performed on smaller local matrices produce better-conditioned continuation operators, which improve numerical stability and reduce sensitivity to round-off errors. Overall, its enhanced robustness and localized error confinement makes it especially well-suited for noisy or non-smooth problems.
@@ 3.4 Error Convergence for Solutions and Derivatives in FC–PINO
For FC Gram, the continuation error on the physical interval consists of two parts, as demonstrated in (Lyon, 2009; Lyon and Bruno, 2010). The dominant term is the interpolation error from matching $d$ boundary values, which converges at order $\mathcal{O}(N^{-d})$. In addition, there is a blending term caused by forcing the continuation to vanish smoothly outside the interval. This blending error does not formally converge with $N$, but can be made numerically negligible. When computing a derivative of order $k$, the interpolation error is reduced to order $\mathcal{O}(N^{-(d-k)})$, while the blending error remains bounded and does not affect the overall convergence rate.
@@
Now, FC–PINO is based on the Fourier neural operator, which enjoys universal approximation for sufficiently smooth operator mappings (Kovachki et al., 2021a), i.e., one can make the approximation error $\|u_{\theta}-\mathcal{G}(f)\|$ arbitrarily small. By the Sobolev embedding theorem, with sufficient regularity, the error in the derivatives can also be controlled and kept arbitrarily small.
@@ 4 Numerical Experiments
We now consider a suite of PDE benchmarks to test the proposed FC–PINO approach. Specifically, we study the 1D self-similar (inviscid) Burgers equation in Section˜4.1, both in the single-instance setting and in the operator-learning setting over a family of parameters, the 2D viscous Burgers equation in Section˜4.2, and the 3D incompressible Navier–Stokes equations for lid-driven cavity flow in velocity–pressure form in Section˜4.3. For each problem, we minimize a physics-informed loss, given by a weighted combination of an interior PDE residual and the appropriate constraint terms.
Derivatives appearing in the physics loss are computed via spectral differentiation in different ways, as discussed in Section˜3.1. To ensure that the reported interior residual losses are comparable and not an artifact of a particular spectral differentiation pipeline, we also validate all of these results using automatic differentiation.
Note that obtaining correct PDE residuals using automatic differentiation can be intricate in practice. Since the FNO outputs a discrete field $u$ on a regular grid rather than a pointwise function, automatic differentiation cannot directly provide derivatives such as $\nabla_{x}u$ at arbitrary coordinates. To enable this verification, we need a differentiable query map $x\mapsto u(x)$ that can be evaluated at continuous points, so that $\nabla_{x}u(x)$ can be obtained by backpropagating through some (or all) of the layers. This can be achieved for instance by replacing the last inverse fast Fourier transform by an inverse discrete Fourier transform, and making sure the resulting computational graphs allow gradients to flow correctly from the query position $x$ to the queried solution $u(x)$ at $x$. Consequently, residual evaluation via automatic differentiation is substantially more memory intensive and computationally expensive than spectral differentiation in this setting, but it provides an important verification of the reported interior residual losses
@@ 4.1.1 Problem Setting
We will train the models using the PDE residual from equation (14) and the boundary condition U(2)=1U(-2)=1. Motivated by the results in (Wang et al., 2022), we introduce an additional smoothness loss term which is the mean squared error of the derivative with respect to $y$ of the left side of equation (14), to improve the smoothness of the solution. Hence the total loss is given by
@@ 4.1.3 Learning the Family of Solutions
It is known that for the Burgers equation to have a smooth self-similar solution, $\lambda$ has to take the form $\lambda=1/(2i+2)$ (Wang et al., 2022). Thus we obtain a collection of values of $\lambda=1/(2i+2)$ by randomly sampling $i\in[0,20]$. By training the neural operator on the corresponding Burgers’ equations, we effectively learn the solution operator over arbitrary values of $\lambda$ corresponding to smooth self-similar solutions.
@@
We first pretrain the proposed FC–PINO architecture with FC–Gram for $3000$ epochs across the selected subset of values of $\lambda$. Then, we fine-tune the pretrained model on individual values of $\lambda$, for $\lambda\in\{1/2,1/12,1/22,1/32,1/52\}$, using the second-order L–BFGS optimizer for 40,00040,\!000 epochs. We embed $\lambda$ in the architecture by representing it as the amplitude of sines and cosines in a NeRF-style embedding (Mildenhall et al., 2021), as
@@
for some integer $K>0$, where $x$ denotes the discretized spatial coordinates.
FNOs naturally leverage sines and cosines through the Fourier transform to learn mappings by capturing interactions between frequency components. Using an encoding with a richer frequency spectrum than a constant $\lambda$ function can allow more efficient learning of complex spectral dynamics. In this numerical experiment, $K=2$ proved sufficient.